Most air traffic passenger use e-tickets. Electronic ticketing allows passengers not worry about paper tickets and it reduces the airlines costs. However in recent times airlines have begun to receive complaints from passengers regarding e-tickets, especially when having to switch planes. To investigate this problem an independent watchdog agency contacted a random sample of 21 airports and collected data on the number of complaints due to e-tickets. The information is as follows;
14.00, 14.00, 16.00, 12.00, 12.00, 14.00, 13.00, 15.00, 16.00, 15.00,
14.00, 12.00, 15.00, 15.00, 14.00, 13.00, 13.00, 12.00, 13.00, 10.00,
13.00
At the .05 level of significance is there evidence to conclude that the ten year average of 15 complaints per month has now changed and actually increased due to electronic ticketing?

Answer by stanbon(75887)

To investigate this problem an independent watchdog agency contacted a random sample of 21 airports and collected data on the number of complaints due to e-tickets. The information is as follows;
14.00, 14.00, 16.00, 12.00, 12.00, 14.00, 13.00, 15.00, 16.00, 15.00,
14.00, 12.00, 15.00, 15.00, 14.00, 13.00, 13.00, 12.00, 13.00, 10.00,
13.00
At the .05 level of significance is there evidence to conclude that the ten year average of 15 complaints per month has now changed and actually increased due to electronic ticketing?
==============
Ho: u = 15
Ha: u > 15
--------
sample mean = 13.57
-----
crit value for one-tail test, alpha=5%, df = 20: t =invT(0.95,20)= 1.7247
-----
test statistic: t(13.57) = (13.57-15)/(1.5/sqrt(21)) = -4.368
------
Conclusion: since the test stat is not in the reject interval, fail
to reject Ho.
======================
Cheers,
Stan H.